Required information

[The following information applies to the questions displayed below.]

Clopack Company manufactures one product that goes through one processing department called Mixing. All raw materials are introduced at the start of work in the Mixing Department. The company uses the weighted-average method of process costing. Its Work in Process T-account for the Mixing Department for June follows (all forthcoming questions pertain to June):

Work in Process—Mixing Department
June 1 balance

30,000

Completed and transferred
to Finished Goods
?
Materials 142,840
Direct labor 91,500
Overhead 109,000
June 30 balance ?

The June 1 work in process inventory consisted of 4,800 units with $17,140 in materials cost and $12,860 in conversion cost. The June 1 work in process inventory was 100% complete with respect to materials and 60% complete with respect to conversion. During June, 37,300 units were started into production. The June 30 work in process inventory consisted of 8,000 units that were 100% complete with respect to materials and 50% complete with respect to conversion.

1 a. Prepare the journal entry to record the overhead cost applied to production. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field.)

b. How many units were completed and transferred to finished goods during the period?

c. Compute the equivalent units of production for materials.

d. Compute the equivalent units of production for conversion.

e. What is the cost of beginning work in process inventory plus the cost added during the period for materials?

f. What is the cost of beginning work in process inventory plus the cost added during the period for conversion?

g What is the cost per equivalent unit for materials? (Round your answer to 2 decimal places.)
Answer #1

1a) Journal entry

date account and explanation Debit Credit
Work in process 109000
Manufacturing overhead 109000
(To record applied overhead)

b) Unit transferred out = 4800+37300-8000 = 34100

c) Equivalent unit of material = 34100+8000 = 42100

d) Equivalent unit of conversion = 34100+(8000*50%) = 38100

e) Cost of beginning work in process plus cost added for material = 17140+142840 = 159980

f) Cost of beginning work in process Plus cost added for conversion = 12860+200500 = 213360

g) Cost per equivalent unit of material = 159980/42100 = 3.80